Train A is leaving Austin to Houston at 82.1 miles per hour.Tran B leaves to Austin at 109.071 miles per hour. How long before they pass eachother on the 240 mile trip ? How far from Austin will they pass ?
Answer by mananth(16946)

Train A is leaving Austin to Houston at 82.1 miles per hour.
Train B leaves to Austin at 109.071 miles per hour.
distance between the two cities = 240 miles..
..
let distance from Austin be x miles
distance from houston = 240-x miles.
...
time = distance / speed
..
when they both meet the time they have travelled is the same.
Train A time to meet = x/82.1
Train B time to meet = 240-x/109.071
..
x/82.1 = 240-x/109.071
109.071x=82.1(240-x)
109.071x= 19704 -82.1x
109.071x+82.1x = 19704
191.171x=19704
x=103.07 miles from austin.
speed of Train A = 82.1
time taken to meet = 103.07/82.1
= 1.256 hours.
They meet after 1.256 hours
...
m.ananth@hotmail.ca